Remember all variances must be designated as Favorable or Unfavorable The Cats Meow manufactures gourmet cat food. Durin
Answer #1

1. Direct material variance =

TUNA = (0.50-0.65)*5500 = $825 unfavorable

Cans = (0.35-0.45)*5000 = 500 unfavorable

2. Direct material quantity variance (tuna) = (5000*1 - 4900)*0.50 = 50 favorable

3. Labor rate variance = (7-8)*5000*0.10 = 500 unfavorable

4. Labor efficiency variance = (5000*0.20 - 5000*0.10)*7 = $3500 favorable